# VideoHallu: Evaluating and Mitigating Multi-modal Hallucinations on Synthetic Video Understanding

### v3

@@ Abstract
Synthetic video generation has gained significant attention for its realism and broad applications, but remains prone to violations of common sense and physical laws. This highlights the need for reliable abnormality detectors that understand such principles and are robust to hallucinations. To address this, we introduce VideoHallu, a benchmark of over 3,000 video QA pairs built from synthetic videos generated by models like Veo2 [1], Sora [2], and Kling [3], paired with expert-crafted counterintuitive QA to evaluate the critical thinking abilities of Multi-modal Large Language Models (MLLMs) on abnormalities that are perceptually obvious to humans but often hallucinated due to language priors. VideoHallu evaluates MLLMs’ abnormality detection abilities with examples across alignment, consistency, commonsense, and physics. We benchmark SOTA MLLMs, including GPT-4o [4], Gemini-2.5-Pro [5], Qwen-2.5-VL [6], Video-R1 [7], and VideoChat-R1 [8]. We observe that these models perform well on many real-world benchmarks like MVBench [9] and MovieChat [10], but still struggle with basic physics-based and commonsense reasoning in synthetic videos. We further show that post-training with Group Relative Policy Optimization (GRPO) [11], using curriculum learning on datasets combining video QA with counterintuitive commonsense and physics reasoning over real and synthetic videos, improves MLLMs’ abnormality detection and critical thinking, demonstrating the value of targeted training for improving their understanding of commonsense and physical laws. Our code is available at https://github.com/zli12321/VideoHallu.git.
@@ 1 Introduction
Figure 1: Challenges in Abnormality Detection in Synthetic Videos. Synthetic videos often contain counterfactual or commonsense-violating contexts misaligned with reality. Prior score-based evaluations focus on video-prompt alignment and visual consistency, overlooking such abnormalities. VideoHallu addresses this gap by collecting synthetic videos with perceptually obvious abnormalities, paired with crafted questions that probe counterintuitive phenomena or test MLLMs’ critical thinking in detecting such abnormalities. SOTA MLLMs evaluated on VideoHallu show severe hallucinations, indicating these models over-rely on language priors and commonsense knowledge rather than ground their answers in the video content.
Video generation has emerged as a pivotal area of research, attracting significant attention due to its potential to revolutionize AI-Generated Content (AIGC) [12, 13], entertainment [14], and robotics [15]. Despite their visual realism, current video generation models [1, 16, 17] often fail to produce visually grounded contexts aligned with real-world commonsense and physical laws, especially when prompted with out-of-distribution descriptions. This stems from the fact that these models are primarily trained to mimic patterns in the data used for training, without an inherent understanding of the physics-based and commonsense principles that govern the real world. As a result, they can generate plausible-looking, but not physically accurate or semantically consistent videos. Meanwhile, current score-based video evaluators, such as VideoScore [18] and DEVIL [19], primarily focus on frame consistency and visual quality. They tend to ignore perceptually obvious violations of the physical laws in videos generated by Runway Gen2 [17] and Sora [16] (Figure 1), providing limited insight into how well synthetic videos align with reality.
This challenge arises from the misalignment between synthetic videos and reality, underscoring the urgent need for explainable evaluation models [20, 21, 22]. Such evaluation approaches are essential not only for systematically identifying unrealistic generations but also for providing interpretable feedback that can guide the improvement of generative models towards producing reality-consistent content. An emerging approach is to leverage Multi-modal Large Language Models (MLLMs) as evaluators of video content, leveraging their commonsense and physics knowledge by breaking down captions or prompts into questions to assess video quality and alignment [23, 24]. Although QA-based evaluation has been effective for real-world videos, it targets visually grounded content that aligns well with the training data of MLLMs, making the questions relatively easy to answer. In contrast, few studies have explored MLLMs’ performance on synthetic videos, a significantly more challenging domain. As shown in Figure 1, SOTA MLLMs like Gemini-2.5-Pro [5], GPT-4o [25], and Qwen2.5-VL [6] often hallucinate when answering counterintuitive questions about synthetic videos. These models rely on prior knowledge, e.g., “a watermelon should break when shot” in the Counterintuitive Question Answering section of Figure 1, rather than the actual video content, where the watermelon implausibly returns intact after exploding in the Sora-generated video. Moreover, MLLMs struggle with critical thinking in synthetic video contexts. They often fail to detect abnormal physical events and instead generate trivial explanations, such as “the bullet missed” or “it was a toy gun,” as seen in Qwen2.5-VL’s response in the Critical Thinking section of Figure 1.
Motivated by such observations, we propose VideoHallu, a benchmark of expert-crafted, diverse QA pairs covering alignment, consistency, commonsense, and physical reasoning to assess MLLMs’ understanding of synthetically generated videos comprehensively. We evaluate several SOTA models and analyze their limitations in detecting synthetic video abnormalities. Furthermore, to enhance their evaluation capabilities on synthetic videos, we apply curriculum learning [26] and Group Relative Policy Optimization (GRPO) [11] to post-train an MLLM using both real-world (VideoLLaVA [27], PhysBench [28]) and synthetic (VideoHallu) datasets. Our goal is to improve MLLMs’ counterintuitive question answering and critical thinking in commonsense and physical reasoning. Our contributions include:
@@
  • We introduce VideoHallu, a novel benchmark for synthetic video understanding and evaluation, comprising over 3,000 expert-annotated QA pairs spanning alignment, spatial-temporal consistency, commonsense, and physical reasoning across 13 sub-categories.
  • We evaluate state-of-the-art video understanding models on VideoHallu. Top-performing MLLMs (e.g., Qwen-2.5-VL (7B/32B), Gemini-2.5-Pro, GPT-4o) achieve only around 50% accuracy, with severe hallucinations in counterintuitive QA and critical thinking tasks.
  • We fine-tune Qwen2.5-VL-7B using GRPO within a curriculum learning pipeline, leveraging a mix of real-world and synthetic video QA sampled from VideoHallu and other video understanding datasets. Our results show that even a small amount of synthetic data (800 samples) leads to a 3% overall accuracy improvement.
2 VideoHallu: Benchmark for Synthetic Video Understanding and Evaluation

### v4

@@ Abstract
Vision–Language Models (VLMs) have achieved remarkable success in video understanding tasks. Yet, a key question remains: do they comprehend visual information, or merely learn superficial mappings between visual and textual patterns? Understanding visual cues, particularly those related to physics and common sense, is crucial for AI systems interacting with the physical world. However, existing VLM evaluations primarily rely on positive-control tests using real-world videos that resemble training distributions. While VLMs perform well on such benchmarks, it is unclear whether they grasp underlying visual and contextual signals or simply exploit visual-language correlations. To fill this gap, we propose incorporating negative-control tests, i.e., videos depicting physically impossible or logically inconsistent scenarios, and evaluating whether models can recognize these violations. True visual understanding should evince comparable performance across both positive and negative tests. Since such content is rare in the real world, we introduce VideoHallu, a synthetic video dataset featuring physics- and commonsense-violating scenes generated using state-of-the-art tools such as Veo2, Sora, and Kling. The dataset includes expert-annotated question–answer pairs spanning four categories of physical and commonsense violations, designed to be straightforward for human reasoning. We evaluate several leading VLMs, including Qwen-2.5-VL, Video-R1, and VideoChat-R1. Despite their strong performance on real-world benchmarks (e.g., MVBench, MMVU), these models hallucinate or fail to detect physical or logical violations, revealing fundamental weaknesses in visual understanding. Finally, we explore reinforcement learning–based post-training on our negative dataset: fine-tuning improves performance on VideoHallu without degrading results on standard benchmarks—indicating enhanced visual reasoning in VLMs. Our data is available at https://github.com/zli12321/VideoHallu.git.
@@ 1 Introduction
Figure 1: Illustrative examples of designed negative-control tests to evaluate the critical thinking abilities of VLMs. Unlike real-world videos, synthetic videos can contain counterfactual or commonsense-violating contexts misaligned with reality. VideoHallu includes such synthetic videos with perceptually obvious abnormalities, paired with crafted questions that probe counterintuitive phenomena or test VLMs’ critical thinking in detecting such abnormalities. When SOTA VLMs are evaluated on VideoHallu, they frequently hallucinate, which suggests that these models rely on language priors and commonsense knowledge rather than truly understand the videos.
Vision–Language Models (VLMs) have made remarkable progress in video understanding. However, they remain prone to hallucinations and shallow visual reasoning [1, 2, 3]. Prior works mitigate these issues across various domains, including chart interpretation [4], video understanding [5], and visual question answering (VQA) [6], primarily through supervised fine-tuning (SFT) or R1-style chain-of-thought training (reinforcement learning) [7, 8]. However, most of these VLM evaluations rely on positive-control test, that is, real-world data drawn from distributions closely aligned with training data. Consequently, it remains unclear whether current VLMs genuinely reason about visual cues or merely exploit prior visual-language correlations within familiar distributions [9].
To truly evaluate visual understanding, mwe test VLMs under negative-control conditions, i.e., videos outside their training distribution that depict physically impossible or logically inconsistent events. These tests reveal whether models detect violations of physics, causality, or commonsense instead of relying on memorized language knowledge. However, constructing such out-of-distribution (OOD) videos in the real world is costly and impractical [10].
Modern video generation models such as Veo2, Sora, and Runway [11, 12, 13, 14] can produce photorealistic but physically impossible scenes. Such models provide an alternative option to generate test videos for probing VLMs’ visual understanding. By careful design, these synthetic videos can be systematically introduced to include violations of gravity, causality, and commonsense interactions [15, 16], enabling controlled OOD evaluations where models need to rely on visual cues. Current VLMs, predominantly trained on videos conforming to physical laws, may thus overfit to statistical regularities rather than learning genuine causal reasoning [17]. Figure 1 illustrates that even state-of-the-art VLMs such as Gemini-2.5-Pro [18], GPT-4o [19], and Qwen2.5-VL [20] hallucinate when confronted with counterintuitive scenarios. For instance, when a watermelon reassembles after an explosion, models rely on linguistic priors (e.g., “a watermelon should break when shot”) rather than actual visual cues, exposing their limited physical reasoning.
To rigorously evaluate such limitations, we introduce VideoHallu, a dataset of expert-curated question–answer pairs over synthetic videos featuring controlled violations of alignment, spatial–temporal consistency, commonsense, and physics. We benchmark several leading VLMs and analyze their failure modes on these physically and logically inconsistent videos. Finally, we explore two post-training strategies, supervised fine-tuning (SFT) and reinforcement learning (RL) via Group Relative Policy Optimization (GRPO) [21], using both real-world data from Video-R1 [7] and synthetic data from VideoHallu. GRPO enhances generalization on synthetic video reasoning without degrading real-world performance.
Contributions.
@@
1)   We introduce VideoHallu, a dataset of 3K expert-annotated QA pairs on synthetic videos that include violations spanning alignment, consistency, commonsense, and physical reasoning.
2)  We evaluate state-of-the-art VLMs and find that even top-performing models (e.g., GPT-4o, Gemini-2.5-Pro) can achieve only $\sim$50% accuracy in our dataset, exhibiting frequent hallucination in counterintuitive scenarios.
3)  GRPO post-training with synthetic data improves visual reasoning on VideoHallu while maintaining real-world performance, providing a path toward more physically grounded VLMs.
2 VideoHallu: Evaluating VLMs’ Synthetic Video Understanding